OpenStudy (anonymous):

How is 3^(3/2) equal to 3sqrt(3) ? steps/explanation? kthxbai

OpenStudy (joannablackwelder):

When you have a fractional exponent, the numerator is the exponent, and the denominator is the root.

OpenStudy (joannablackwelder):

So, we can rewrite like this:

OpenStudy (joannablackwelder):

Make sense so far, @ProfessionalDEEK?

OpenStudy (anonymous):

\[\sqrt{3^3}=\sqrt{3^2*3}=3\sqrt{3}\]

OpenStudy (michele_laino):

please note that: \[3\sqrt{3}=3\left( 3 \right)^{1/2}=3^{1+1/2}=...\] what is \[1+\frac{ 1 }{ 2 }=...\]